# Sharp one-point estimates and Minkowski content for the scaling limit of three-dimensional loop-erased random walk

### v1

@@ 1.2 Proof overview
Acknowledgements: SHT would like to acknowledge the support of UNAM-PAPIIT grant IA103724. XL is supported by National Key R&D Program of China (No. 2021YFA1002700 and No. 2020YFA0712900) and NSFC (No. 12071012). DS was supported by JSPS Grant-in-Aid for Scientific Research (C) 22K03336, JSPS Grant-in-Aid for Scientific Research (B) 22H01128 and 21H00989.
@@ 2.1 General notation
We call an infinite sequence of points $\lambda=[\lambda(0),\lambda(1),\ldots]$ an infinite path, if $|\lambda(j+1)-\lambda(j)|=m^{-1}$ for all $j=0,1,\ldots$. We call $\lambda$ transient if for any $a\in m^{-1}{\mathbb{Z}}^{3}$, |$|\{k:\lambda(k)=a\}|<\infty$. For a transient infinite path, we can define its loop erasure, denoted as ${\rm LE}(\lambda)$, in a manner similar to Definition 2.1.
We write ||$\cdot$|| ⋅ | for the Euclidean distance in ${\mathbb{R}^{3}}$. Given $B_{1},\ldots,B_{n}$ a sequence of Borel subsets of ${\mathbb{R}^{3}}$, let
@@
$B(G,r)\coloneqq\{x\in{\mathbb{R}^{3}},{\rm dist}(x,G)\leq r\}$ (2.2)
@@ 2.3 Minkowski content
for its $\delta$-dimensional Minkowski content, provided the limit exists. Note that for each $A$, there exists at most one $\delta$ such that its $\delta$-dimensional content is finite and positive. The definition of the Minkowski content depends on the volume. Then, when the corresponding limit exists, the Minkowski content is scale-invariant and invariant under the isometries of ${\mathbb{R}}^{d}$.
@@
To this end, we employ a classical method, which involves proving the existence of one-point and two-point Green’s functions with some convergence rate. The respective conditions and claims are summarized in the following proposition (quoted from [8, Proposition 4.8], with some adaptations, see Remark 2.3).
@@
[…]
@@
Compared to [8, Proposition 4.8], in the proposition above, we have weakened the conditions on the convergence rates of one-point and two-point functions in (2.7). More precisely, the version in [8] is equivalent to taking $b=1$ here. Note that this modification does not change the argument. More precisely, the RHS of (4.41) and (4.42), ibid., stay summable even when $e^{-us}$ is replaced by $e^{-us^{b}}$ for some $b>0$, which implies the RHS of (4.43) and (4.44), ibid., still converges to 00. We have made this modification to suit the convergence rate we have obtained in [20], as stated in Proposition 2.14, in particular the error terms in (2.33) and (2.34).
@@
By the same argument as in [8, Appendix A], one arrives at the following proposition.
@@
indeed induces a measure. However, for random fractals such as the scaling limit of LERW, one can check that conditions of Proposition 2.2 hold for boxes that are not too close to the boundary of the domain or the starting point of the curve). This justifies our choice of ${\cal D}^{o}$.
@@ 2.4 Weak convergence of probability measures
In this subsection, we let ${\cal M}(\overline{\mathbb{D}})$ be the space of all finite measures on $\overline{\mathbb{D}}$ equipped with the topology of the weak convergence. The space ${\cal M}(\overline{\mathbb{D}})$ is complete, metrizable and separable (see Section 6, in particular [3, Theorem 6.8]).
We now briefly recall basic facts on the weak convergence of probability measures here. See [4, Chapter 3] for details and proofs.
@@ 2.5 The scaling limit of 3D LERW
It is noteworthy that the scaling invariance of ${\cal K}$ actually implies that the convergence in (2.18) follows not only along dyadic scales, but also as $m\to\infty$ continuously. See [20, Subsection 3.7] for a detailed explanation of the proof of the following proposition.
@@
$\eta_{m}\overset{\rm w}{\to}{\cal K}\mbox{ as $m\to\infty$}$
@@
$\Gamma\coloneqq\big{\{}\gamma\in{\cal H}(\overline{\mathbb{D}})\ |\ \gamma\text{ is a simple curve satisfying }\gamma^{s}=0\text{ and }\gamma\cap\partial\mathbb{D}=\{\gamma^{e}\}\big{\}}.$ (2.22)
@@
In [20], the last two authors improve the convergence result by showing that the (renormalized) occupation measure of LERW has a scaling limit and rescaled LERW converges as a naturally parametrized curve to the curve given through parametrizing ${\cal K}$ by the limiting occupation measure, as discussed in Subsection 1.1. More precisely, we have the following theorem. Recall
@@
be the occupation measure of LERW scaled by the one-point function.
@@
As $n\in{\mathbb{Z}^{+}}$ tends to infinity,
$({\cal K}_{2^{n}},\mu_{2^{n}})\xrightarrow{\rm w}({\cal K},\mu)$ (2.23)
@@
The one-point estimates over dyadic scales along dyadic scales in [19], allow taking the scaling factor 1.2 as $f_{2^{n}}=2^{\beta n}$. Recall that we write
@@
and $\overline{\eta}_{2^{n}}$ is the LERW on $\mathbb{D}_{2^{n}}$ parametrized by $\overline{\mu}_{2^{n}}$ (compare (1.16)). Therefore, the results in [19, 20] imply that $\bar{\eta}_{2^{n}}$ converges to $\eta$ in natural parametrization. In other words, as $n\in{\mathbb{Z}^{+}}$ tends to infinity,
$\eta_{2^{n}}\xrightarrow{\rm w}\eta$
@@ 2.6 Up-to-constant one- and two-point estimates
The following estimates follow from [20, Equation (4.20)] and [20, Corollary 6.4].
@@
For $1000<r<d_{x}/$ 100, there exists some $M_{0}(x,r)$ such that for all $m\geq M_{0}$,
@@
In [19], the last two authors obtain the following sharp one-point function asymptotics (see Theorem 1.1, ibid.). Recall that $\eta^{m}$ stand for the LERW on $\mathbb{D}_{m}$ as defined in Section 1.
@@
We now turn to two-point estimates. Recall the definition of ${\cal D}^{o}$ in (2.15). Let $V\in{\cal D}^{o}$. Write $x_{V}$ for the center of $V$ and $C_{V}=C_{x_{V}}$ where $C_{x_{V}}$ is defined in (2.24). For two different points $z,w\in V$, write $d=|z-w|$.
@@
For $1000<r,r^{\prime}<d/$ 100 such that $B(z,r),B(w,r^{\prime})\subset V$, there exists some $M_{1}(V,d,r,r^{\prime})>0$ such that for all $m\geq M_{1}$, we have
@@
We omit the proof as it is essentially the same as those of Propositions 4.2 and 4.3 of [20] (see also [20, Remark 4.5]). Note that the independence of $M_{1}$ on $z$ and $w$ follows from the equivalence of the convergence under Lévy metric and the weak convergence under $d_{\text{Haus}}$-metric.
@@ 2.7 Asymptotic independence for 3D LERW
For $k\in{\mathbb{Z}}^{+}$, consider $m\in\mathbb{R}^{+}$ such that $m\geq 2^{k^{4}}$. Let $z,w\in V\cap m^{-1}{\mathbb{Z}}^{3}$ such that $|z-w|\geq 2^{-k^{2}}$. Let
@@
This proposition follows from a minor modification of the proof for (5.11), (5.6) and (5.7) of [20] where $S_{1}$ and $S_{2}$ are restricted to certain boxes. See also (5.8) and (5.9), ibid. Note that this decoupling result holds for non-dyadic scales too.
@@ 3 A continuity result
Let $\eta=\eta_{2^{n}}$ be the LERW on $D_{2^{n}}$ from the origin. For $a>b\geq 10^{6}$, we set
@@
$F^{i}_{+}=\big{\{}(y_{1},y_{2},y_{3})\in\partial B\ \big{|}\ y_{i}\geq x_{i}\big{\}}\ \ \ \ \ \text{ and }\ \ \ \ \ F^{i}_{-}=\big{\{}(y_{1},y_{2},y_{3})\in\partial B\ \big{|}\ y_{i}\leq x_{i}\big{\}}.$
Note that $\partial B=\bigcup_{i=1}^{3}F^{i}_{+}\cup F^{i}_{-}$. Let $\tau$ be the first time that $\gamma$ hits $B$. Then we have
@@
We now bound the RHS of (3.1). By symmetry we can deal with the case of $F=F^{3}_{-}$ only. Let
=$\ell=\big{\{}y=(y_{1},y_{2},y_{3})\in\partial B\ \big{|}\ (y_{1},y_{2})=(x_{1},x_{2})\big{\}}$
@@
By (2.25), $P\big{(}\eta\cap B(y_{j},2\zeta)\neq\emptyset\big{)}\leq C\zeta^{3-\beta}$ for each $j$. Since $02-\beta>$ 0, we see that
@@
Once we have (3.4), we can follow the same scenario as in the proof of [20, Proposition 6.2]. The steps are as follows.
@@
    The coupling argument via a cut time to prove that [20, Equation (5.10)] still works here. The same argument shows that there exists $\delta>0$ such that
@@
    =\;= $P\big{(}\eta\cap B\neq\emptyset,\ \eta(\tau)\in F\big{)}-P\big{(}\eta\cap B\neq\emptyset,\ \eta(\tau)\in F,\ \eta\cap\widehat{B}_{a}=\emptyset\big{)}$
[…]
    $\ \ \ \ \ -P\big{(}\eta\cap B\neq\emptyset,\ \eta(\tau)\in F,\ \eta\cap\widehat{B}_{a}=\emptyset\big{)}.$
Therefore, it follows from (3.6) that
@@
[…]
@@
for some $c,C\in(0,\infty)$. Since we can couple $\eta=\eta_{2^{n}}$ and ${\cal K}$ in the same probability space such that the Hausdorff distance between them converges to 00 as $n\to\infty$, it follows from (3.7) that
@@ 4 Sharp one-point function estimates
  • (Initial value) there exist some $A$ and $B$ such that
@@
  • (Local logarithmic additivity) For all $r,s\in[0,1]$,
@@
We omit the proof as it follows from classical methods of solving functional equations.
We now turn back to one-point functions. For the rest of this subsection, fix $x\in\mathbb{D}\setminus\{0\}$ and for any $r\in{\mathbb{R}^{+}}$, write
$z(r)=P(x_{2^{r}}\in\eta_{2^{r}}).$
@@
Let $G_{\mathbb{D}}(\cdot,\cdot)$ stand for the Green’s function of the simple random walk in $\mathbb{D}$ killed upon exiting $\mathbb{D}$. Given $M>0$, for $x,y\in\mathbb{D}_{M}$, let $P^{M}_{x,y}$ stand for the joint law of $(X,Y)$ where $X$ is a simple random walk started from $x_{M}$ conditioned to hit 00 before exiting $\mathbb{D}_{M}$ and $Y$ is a simple random walk started from $y_{M}$ stopped at exiting $\mathbb{D}_{M}$. As in [19, Lemma 5.1], the decomposition of the non-intersection probability in [25, Proposition 8.1] and the reversibility of LERW give that
$z({M})=G_{\mathbb{D}_{M}}(0,x_{M})P^{M}_{x_{M},x_{M}}({\rm LE}(X)\cap Y[1,T]=\emptyset).$
By the same technique as in [19], we need to change the starting point of $X$’s and $Y$’s as well as truncate the tail of $X$. Let $m=2^{n}$ and
$x_{1}=(2^{-5qn-r-s},0,0),\;y_{1}=-x_{1},\;x_{2}=sx_{1},\;y_{2}=sy_{1}.$
For $x_{*},y_{*}\in\mathbb{R}^{3}$, let $x^{M}=(x-x_{*})_{2^{M}}$ and $y^{M}=(x+y_{*})_{2^{M}}$ and write
$W_{x_{*},y_{*}}(M)\coloneqq P^{2^{M}}_{x^{M},y^{M}}\big{(}{\rm LE}(X)\cap Y=\emptyset\big{)}$
for the non-intersection probability of the loop-erasure of a SRW started from $x^{M}$ conditioned to hit 00 before exiting $\mathbb{D}_{2^{M}}$ and SRW $Y$ starting from (the discretization of) $y^{M}$ and stopped at exiting $\mathbb{D}_{2^{M}}$.
@@
We start with (1.7). We first note that in [19, Theorem 1.1] the claim (1.7) has already been verified for $m=2^{n}$, $n\in{\mathbb{Z}^{+}}$. Moreover, the proof of [19, Theorem 1.1] is valid for $m\in[2^{n},2^{n+1})$, with $n\in{\mathbb{Z}^{+}}$. We write $m=2^{n+r}$ for $r\in[0,1)$, and following the steps in [19], by an application of [10] for the scaling limit of LERW in $2^{s}\mathbb{D}$ for $s\in[0,1]$ [^2] Note that for one can pick constants in that proof uniformly for $s\in[0,1]$, for the same reason as the proof of Proposition 2.6., we obtain that there exist $c,\delta>0$ and some $f:[0,1)\to\mathbb{R}^{+}$, such that for any $r\in[0,1)$,
$z(m)=f(r)2^{-(3-\beta)n}\Big{[}1+O\big{(}d_{x}^{-c}2^{-\delta n}\big{)}\Big{]}$ (4.8)
@@
Now we verify that the function $f$ satisfies the conditions of Proposition 4.1. The claim (4.1) is trivial. Note that condition (4.2) follows from Proposition 4.2, as the constants in the $O(\cdot)$ expression in (4.8) does not depend on $r$. To check condition (4.3), we apply [25, Proposition 8.2], [25, Proposition 8.5] and [19, Corollary 1.3]. This finishes the proof of (1.7).
Recall the definition of $g(\cdot)$ from (1.7). The continuity of $g(\cdot)$ follows from [20, Lemma 3.3]. The rotation-invariance of $g(\cdot)$ is a corollary of Theorem 1.4 and the rotation-invariance of ${\cal K}$. ∎
Remark 4.3.
Although open at the moment, we conjecture that $g(\cdot)$ is differentiable and satisfies some asymptotics which improves (1.5) and (1.6) in [19].
Finally, we note that similar techniques as in the proof above also allow us to obtain the sharp asymptotics of non-intersection probabilities of LERW and SRW in any scale, which is an improvement of [19, Theorem 1.2].
Let $S=S[0,\infty)$ be a discrete-time simple random walk on $\mathbb{Z}^{3}$ started from 00 and $T_{m}\coloneqq\inf\{T\geq 0;|S[T]|\geq m\}$ be the first exit time of a ball of radius $m$. Let $\gamma_{m}\coloneqq{\rm LE}(S[0,T_{m}])$ be the LERW of SRW truncated upon leaving a ball of radius $m$. Recall that $\beta\in(1,5/3]$ stands for the growth exponent of 3D LERW. Let $S^{\prime}$ be an independent copy of $S$ and define $T^{\prime}_{m}$ accordingly. Write $P$ for the joint law of $S$ and $S^{\prime}$. Let
${\rm Es}(m)\coloneqq P\Big{(}\gamma_{m}\cap S^{\prime}[0,T^{\prime}_{m}]=\emptyset\Big{)}$ (4.11)
stand for the non-intersection probability of $\gamma_{m}$ and $S^{\prime}[0,T^{\prime}_{m}]$, and recall the definition of the universal constant $c>0$ from [19, Theorem 1.2].

### v2

@@ 2.1 General notation
We call an infinite sequence of points $\lambda=[\lambda(0),\lambda(1),\ldots]$ in $m^{-1}{\mathbb{Z}}^{3}$ an infinite path, if $|\lambda(j+1)-\lambda(j)|=m^{-1}$ for all $j=0,1,\ldots$. We call $\lambda$ transient if for any $a\in m^{-1}{\mathbb{Z}}^{3}$, $|\{k:\lambda(k)=a\}|<\infty$. For a transient infinite path, we can define its loop erasure, denoted as ${\rm LE}(\lambda)$, in a manner similar to Definition 2.1.
We write $|\cdot|$ for the Euclidean distance in $\mathbb{R}^{3}$}. Given $B_{1},\ldots,B_{n}$ a sequence of Borel subsets of $\mathbb{R}^{3}$}, let
@@
$B(G,r)\coloneqq\{x\in{\mathbb{R}^{3}}\,|\,{\rm dist}(x,G)\leq r\}$ (2.2)
@@ 2.3 Minkowski content
for its $\delta$-dimensional Minkowski content, provided the limit exists. Note that for each $A$, there exists at most one $\delta$ such that its $\delta$-dimensional content is finite and positive. The definition of the Minkowski content depends on the volume. Then, when the corresponding limit exists, the Minkowski content enjoys a scaling property and is invariant under the isometries of ${\mathbb{R}}^{d}$.
@@
To this end, we employ a classical method, which involves proving the existence of one-point and two-point Green’s functions with some convergence rate. The respective conditions and claims are summarized in the following proposition (quoted from [9, Proposition 4.8], with some adaptations, see Remark 2.3).
@@
$J_{V}={\rm Cont}_{\delta}(V\cap A),\mbox{ a.s.},$ (2.9)
$\mu\left[J_{V}\right]=\lim_{s\rightarrow\infty}\mu\left[J_{s,V}\right]=\int_{V}G(z)\,dz,\mbox{ and}$ (2.10)
$\mu\left[J_{V}^{2}\right]=\lim_{s\rightarrow\infty}\mu\left[J_{s,V}^{2}\right]=\int_{V}\int_{V}G(z,w)\,dz\,dw.$ (2.11)
@@
Compared to [9, Proposition 4.8], in the proposition above, we have weakened the conditions on the convergence rates of one-point and two-point functions in (2.7). More precisely, the version in [9] is equivalent to taking $b=1$ here. Note that this modification does not change the argument. More precisely, the RHS of (4.41) and (4.42), ibid., stay summable even when $e^{-us}$ is replaced by $e^{-us^{b}}$ for some $b>0$, which implies the RHS of (4.43) and (4.44), ibid., still converges to 0. We have made this modification to suit the convergence rate obtained in [22], as stated in Proposition 2.14, in particular the error terms in (2.33) and (2.34).
@@
By the same argument as in [9, Appendix A], one arrives at the following proposition.
@@
indeed induces a measure. However, for random fractals such as the scaling limit of LERW, one can check that conditions of Proposition 2.2 hold for boxes that are not too close to the boundary of the domain or the starting point of the curve. This justifies our choice of ${\cal D}^{o}$.
@@ 2.4 Weak convergence of probability measures
In this subsection, we let ${\cal M}(\overline{\mathbb{D}})$ be the space of all finite measures on $\overline{\mathbb{D}}$ equipped with the topology of the weak convergence. The space ${\cal M}(\overline{\mathbb{D}})$ is complete, metrizable and separable (see Section 6, in particular [4, Theorem 6.8]).
We now briefly recall basic facts on the weak convergence of probability measures here. See [5, Chapter 3] for details and proofs.
@@ 2.5 The scaling limit of 3D LERW
It is noteworthy that the scaling invariance of $\cal K$} actually implies that the convergence in (2.18) follows not only along dyadic scales, but also as $m\to\infty$ continuously. See [22, Subsection 4.1] for a detailed explanation of the proof of the following proposition.
@@
$\eta_{m}\overset{\rm w}{\to}{\cal K}\qquad\mbox{ as }\qquad m\to\infty$
@@
[…]
@@
In [22], the last two authors improve the convergence result by showing that the (renormalized) occupation measure of LERW has a scaling limit and rescaled LERW converges as a naturally parametrized curve to the curve given through parametrizing $\cal K$} by the limiting occupation measure, as discussed in Subsection 1.1. More precisely, we have the following theorem. Recall that
@@
is the occupation measure of LERW scaled by the one-point function.
@@
As $m\in{\mathbb{R}^{+}}$ tends to infinity,
$({\cal K}_{m},\mu_{m})\xrightarrow{\rm w}({\cal K},\mu)$ (2.23)
@@
The one-point estimates over dyadic scales along dyadic scales in [20], allow taking the scaling factor $f_{2^{n}}$ in (1.2) as a multiple of $2^{\beta n}$, for $n\in\mathbb{Z}^{+}$. Recall that we write
@@
and $\overline{\eta}_{2^{n}}$ is the LERW on $\mathbb{D}_{2^{n}}$ parametrized by $\overline{\mu}_{2^{n}}$ (compare (1.16)). Therefore, the results in [20, 22] imply that $\bar{\eta}_{2^{n}}$ converges weakly to some $\overline{\eta}$ in natural parametrization. In other words, as $n\in{\mathbb{Z}^{+}}$ tends to infinity,
$\overline{\eta}_{2^{n}}\xrightarrow{\rm w}\overline{\eta}$
@@ 2.6 Up-to-constant one- and two-point estimates
The following estimates follow from [22, Equation (5.13)] and [22, Corollary 7.4].
@@
For $0<r<d_{x}/100$, there exists some $M_{0}(x,r)$ such that for all $m\geq M_{0}$,
@@
In [20], the last two authors obtain the following sharp one-point function asymptotics (see Theorem 1.1, ibid.) along dyadic scales. Recall that $\eta^{m}$ stand for the LERW on $\mathbb{D}_{m}$ as defined in Section 1.
@@
We now turn to two-point estimates. Recall the definition of ${\cal D}^{o}$ in (2.15). Let $V\in{\cal D}^{o}$. Write $x_{V}$ for the center of $V$ and $C_{V}=C_{x_{V}}$ where $C_{x_{V}}$ is defined in (2.24). For two different points $z,w\in V$, write d=|zw|d=|z-w|.
@@
For $0<r,r^{\prime}<d/100$ such that $B(z,r),B(w,r^{\prime})\subset V$, there exists some $M_{1}(V,d,r,r^{\prime})>0$ such that for all $m\geq M_{1}$, we have
@@
We omit the proof as it is essentially the same as those of Propositions 5.2 and 5.3 of [22]. Note that the independence of $M_{1}$ on $z$ and $w$ follows from the equivalence of the convergence under Lévy metric and the weak convergence under $d_{\text{Haus}}$-metric.
@@ 2.7 Asymptotic independence for 3D LERW
For $k\in{\mathbb{Z}}^{+}$, consider $m\in\mathbb{R}^{+}$ such that $m\geq 2^{k^{4}}$ and
for any $S\subset m^{-1}\mathbb{Z}^{3}$. Let $z,w\in V\cap m^{-1}{\mathbb{Z}}^{3}$ such that $|z-w|\geq 2^{-k^{2}}$. Let
@@
This proposition follows from a minor modification of the proof for (6.6) and (6.7) of [22] where $S_{1}$ and $S_{2}$ are restricted to certain boxes. See also (6.8) and (6.9), ibid. Note that this decoupling result holds for non-dyadic scales too.
@@ 3 A continuity result
Figure 1: The equator $\ell$ divides $\partial B$ into two shells: the northern hemisphere $F_{+}^{3}$ and the southern one $F=F_{-}^{3}$. The shaded area represents the set $F^{\prime}\subset\partial F$, consisting of points $x\in F$ in the southern hemisphere that are at a distance greater than $\zeta=\varepsilon^{b/20}$ from the equatorial line $\ell$.
@@
Let $\eta=\eta_{2^{n}}$ be the LERW on $\mathbb{D}_{2^{n}}$ from the origin. For $a>b\geq 10^{6}$, we set
@@
[…]
Note that $\partial B=\bigcup_{i=1}^{3}F^{i}_{+}\cup F^{i}_{-}$. Let $\tau$ be the first time that $\eta$ hits $B$. Then we have
@@
We now bound the RHS of (3.1). By symmetry, we can deal with the case of $F=F^{3}_{-}$ only. See Figure 1 for reference. Let
$\ell=\big\{y=(y_{1},y_{2},y_{3})\in\partial B\ \big|y_{3}=x_{3}\big\}$
@@
By (2.25), $P\big(\eta\cap B(y_{j},2\zeta)\neq\emptyset\big)\leq C\zeta^{3-\beta}$ for each $j$. Since $2-\beta>0$, we see that
@@
Once we have (3.4), we can follow the same scenario as in the proof of [22, Proposition 7.2]. The steps are as follows.
@@
    The coupling argument via a cut time to prove [22, Equation (7.10)] still works here. The same argument shows that there exists $\delta>0$ such that
@@
    $=\;$ $P\big(\eta\cap B\neq\emptyset,\ \eta(\tau)\in F\big)-P\big(\eta\cap B\neq\emptyset,\ \eta(\tau)\in F,\ \eta\cap\widehat{B}_{a}=\emptyset\big)$
    $\leq\;$ $P\big(\widetilde{\eta}\cap\widehat{B}_{a}\neq\emptyset,\ \widetilde{\eta}(\widetilde{\tau})\in F\big)+Cr\zeta^{2-\beta}-P\big(\eta\cap B\neq\emptyset,\ \eta(\tau)\in F,\ \eta\cap\widehat{B}_{a}=\emptyset\big).$
It follows from (3.6) that
@@
$\leq\;$ $P\big(\widetilde{\eta}\cap\widehat{B}_{a}\neq\emptyset,\ \widetilde{\eta}(\widetilde{\tau})\in F\big)-P\big(\eta\cap\widehat{B}_{a}\neq\emptyset,\ \eta(\tau)\in F\big)+Cr\zeta^{2-\beta}$
$\leq\;$ $C\varepsilon^{\delta b}+Cr\zeta^{2-\beta}.$
@@
for some $c,C\in(0,\infty)$. Since we can couple $\eta=\eta_{2^{n}}$ and $\cal K$} in the same probability space such that the Hausdorff distance between them converges to 0 as $n\to\infty$, it follows from (3.7) that
@@ 4 Sharp one-point function estimates
  • (Boundary value) there exist some $A$ and $B$ such that
@@
  • (Local logarithmic additivity) For all $r,s\geq 0$ satisfying $\lfloor r\rfloor=\lfloor s\rfloor$,
@@
Proof.
By an induction argument, it is straightforward to check that (4.1) and (4.2) imply that (4.4) holds for all integers of the form $r=2^{n}>0$. If for some $r_{0}\in(0,1]$ the identity (4.4) does not hold, then by iterating (4.2) one can find a large integer $2^{M}$ such that $f(2^{M}r_{0})/f(\lfloor 2^{M}r_{0}\rfloor)\notin(1/C,C)$, contradicting (4.3)! It follows that (4.4) holds for all $r\in(0,1]$.
From here, we extend (4.4) to all positive $r=\cup_{m\in\mathbb{N}}(0,m]$ by induction on $m$. Assuming that (4.4) holds for all $r\in(0,m]$, let $t\in(m,m+1]$. Applying (4.2) to the decomposition $t=t/2+t/2$, we get that $t$ also satisfies (4.4). ∎
We now turn back to one-point functions.
Our first tool is a decomposition of the one-point probability into a Green’s function and a certain non-intersection probability. Let $X$ be a simple random walk started from $x_{m}$ conditioned to hit 0 before exiting $\mathbb{D}_{m}$, and let $Y$ be a simple random walk started from $y_{m}$ stopped at exiting $\mathbb{D}_{m}$ (we denote this stopping time by $T$). We write $P_{x,y}^{m}$ for the joint law of $(X,Y)$. (Note that we omit the scaling for the starting points $x_{m}$ and $y_{m}$, to lighten the notation). From [20, Lemma 5.1] we have that
$P(x_{m}\in\eta_{m})=G_{\mathbb{D}_{m}}(0,x_{m})P^{m}_{x,x}({\rm LE}(X)\cap Y[1,T]=\emptyset).$ (4.5)
This last identity follows from [20, Lemma 5.1]. A version of this decomposition was proved for homogeneous Markov chains in [2, Proposition 5.2].
For the rest of this subsection, fix $x\in\mathbb{D}\setminus\{0\}$ and for any $r\in{\mathbb{R}^{+}}$, write
$z(r)=P(x_{2^{r}}\in\eta_{2^{r}}).$ (4.6)
@@
Let $G_{\mathbb{D}}(\cdot,\cdot)$ stand for the Green’s function of the simple random walk in $\mathbb{D}$ killed upon exiting $\mathbb{D}$. Given $r>0$ and $x,y\in\mathbb{D}_{2^{r}}$, recall that $P^{2^{r}}_{x,y}$ denotes the joint law of $(X,Y)$ defined just above (4.5). As in (4.5), the probability in (4.6) that $x_{2^{r}}$ is on $\eta_{2^{r}}$ satisfies
$z({r})=G_{\mathbb{D}_{2^{r}}}(0,x_{2^{r}})P^{2^{r}}_{x,x}({\rm LE}(X)\cap Y[1,T]=\emptyset).$ (4.8)
By the same technique as in [20], we need to change the starting point of $X$’s and $Y$’s as well as truncate the tail of $X$. We pick a small $q\in(0,1/10)$ as in [20, Proposition 3.4]. For any $n\in\mathbb{Z}^{+}$ and $r,s\in[0,1]$, let
$x_{1}=(2^{-5qn-r-s},0,0),\;y_{1}=-x_{1},\;x_{2}=2^{s}x_{1},\;y_{2}=2^{s}y_{1}.$
For $x_{i},y_{i}\in\mathbb{R}^{3}$, $i=1,2$, write
$W_{x_{i},y_{i}}(r)\coloneqq P^{2^{r}}_{(x-x_{i}),(x+y_{i})}\big({\rm LE}(X)\cap Y=\emptyset\big)$
for the non-intersection probability of the loop-erasure of a SRW started from (the discretization of) $x-x_{i}$ conditioned to hit 0 before exiting $\mathbb{D}_{2^{r}}$ and SRW $Y$ starting from (the discretization of) $x+y_{i}$ and stopped at exiting $\mathbb{D}_{2^{r}}$.
@@
We start with (1.7). We first note that in [20, Theorem 1.1] the claim (1.7) has already been verified for $m=2^{n}$, $n\in{\mathbb{Z}^{+}}$. Moreover, the proof of [20, Theorem 1.1] is valid for $m\in[2^{n},2^{n+1})$, with $n\in{\mathbb{Z}^{+}}$. We write $m=2^{n+r}$ for $r\in[0,1)$, and following the steps in [20], by an application [^2] Note that one can pick constants in that proof uniformly for $s\in[0,1]$, for the same reason as the proof of Proposition 2.6. of [11] for the scaling limit of LERW in $2^{s}\mathbb{D}$ for $s\in[0,1]$, we obtain that there exist $c,\delta>0$ and some $f:[0,1)\to\mathbb{R}^{+}$, such that for any $r\in[0,1)$,
$z(n+r)=f(r)2^{-(3-\beta)n}\Big[1+O\big(d_{x}^{-c}2^{-\delta n}\big)\Big]$ (4.11)
@@
Now we verify that the function $f$ satisfies the conditions of Proposition 4.1. The claim (4.1) is trivial. Note that condition (4.2) follows from Proposition 4.2, as the constants in the $O(\cdot)$ expression in (4.11) does not depend on $r$. To check condition (4.3), we apply [27, Proposition 8.2], [27, Proposition 8.5] and [20, Corollary 1.3]. This finishes the proof of (1.7).
Recall the definition of $g(\cdot)$ from (1.7). The continuity of $g(\cdot)$ follows from [22, Lemma 5.4]. The rotation-invariance of $g(\cdot)$ is a corollary of Theorem 1.4 and the rotation-invariance of $\cal K$}. ∎
Finally, we note that similar techniques as in the proof above also allow us to obtain the sharp asymptotics of non-intersection probabilities of LERW and SRW in any scale, which is an improvement of [20, Theorem 1.2].
Let $S=S[0,\infty)$ be a discrete-time simple random walk on $\mathbb{Z}^{3}$ started from 0 and $T_{m}\coloneqq\inf\{T\geq 0;|S(T)|\geq m\}$ be the first exit time of a ball of radius $m$. Let $\gamma_{m}\coloneqq{\rm LE}(S[0,T_{m}])$ be the LERW of SRW truncated upon leaving a ball of radius $m$. Recall that $\beta\in(1,5/3]$ stands for the growth exponent of 3D LERW. Let $S^{\prime}$ be an independent copy of $S$ and define $T^{\prime}_{m}$ accordingly. Write $P$ for the joint law of $S$ and $S^{\prime}$. Let
${\rm Es}(m)\coloneqq P\Big(\gamma_{m}\cap S^{\prime}[1,T^{\prime}_{m}]=\emptyset\Big)$ (4.14)
stand for the non-intersection probability of $\gamma_{m}$ and $S^{\prime}[1,T^{\prime}_{m}]$ and recall the definition of the universal constant $c>0$ from [20, Theorem 1.2].
@@
The sharp estimate on the non-intersection probability allows us to improve the upper bound on the one-ball hitting probability given in [26, Lemma 7.1]. We state this improved bound in the following lemma, as it will be useful to show our sharp bound for the one-ball hitting probability.
Lemma 4.4.
There exists a universal constant $C<\infty$ such that for all $m\in\mathbb{R}^{+}$, $s\in(0,1)$ and each $x\in\mathbb{D}_{m}$ with $d_{x}>2s$, we have
$P\left(\eta_{m}\cap B(x,s)\neq\emptyset\right)\leq C\left(\frac{s}{|x|}\right)^{3-\beta}.$
The proof of Lemma 4.4 is the same as that of [26, Lemma 7.1], but with the bounds for the escape probability from [27, Lemma 7.2.2] replaced by those in Theorem 4.3.